the function y=20*.975^x models the intensity of sunlight beneath the surface of the ocean. the output y represents the percent of surface sunlight intensity that reachs a depth of x feer. the model is accurate from about 20 feet to about 600 feet beneath the surface. find the percent of sunlight 50 feet beneath the surface. find the percent of sunlight at a depth of 370 feet.
Answer by stanbon(57282)

the function y=20*.975^x models the intensity of sunlight beneath the surface of the ocean.
----------------------------
the output y represents the percent of surface sunlight intensity that reaches a depth of x feet.
-------------------------------
the model is accurate from about 20 feet to about 600 feet beneath the surface. find the percent of sunlight 50 feet beneath the surface.
y=20*.975^x
---
y = 20*0.975^50
y = 20*0.2820
y = 5.64%
-------------------

find the percent of sunlight at a depth of 370 feet.
y = 20*.975^x
y = 20*0.975^370
y = 20*0.00008545
y = 0.0017%
=================
Cheers,
Stan H.
=================